OpenStudy (anonymous):

Please Help. The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2).

OpenStudy (anonymous):

Answer choices y = −2x + 4 y = negative 1 over 2x − 1 over 2 y = 1 over 2x − 7 over 2 y = 2x − 8

OpenStudy (anonymous):

Parallel lines have the same slope, from y=mx+b m values must be equal and b is the value of the y-intercept